OpenStudy (anonymous):

Calculate the Molarity of a) .250moles of Na2S in 1.70 L of solution b) 32.7grams of MgS in 857 mL of solution Please help! medals given

OpenStudy (cggurumanjunath):

http://www.occc.edu/kmbailey/Chem1115Tutorials/Molarity.htm

OpenStudy (cggurumanjunath):

go through this link @Loveforhelp .

OpenStudy (cggurumanjunath):

Molar mass of Na2S = 78.04454 g/mol

OpenStudy (cggurumanjunath):

Percent composition by element Element Symbol Atomic Mass # of Atoms Mass Percent Sodium Na 22.989770 2 58.914% Sulfur S 32.065 1 41.086%

OpenStudy (anonymous):

ok i am thank you.

OpenStudy (cggurumanjunath):

@Loveforhelp

OpenStudy (anonymous):

for a. i got .14 for b. i got .677

OpenStudy (anonymous):

first one wrong. :/ @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

first one wrong. :/ ?

OpenStudy (cggurumanjunath):

can u show me ur steps !

OpenStudy (anonymous):

for A. i did mols of solute/L of solution.... = .250mols/1.70L

OpenStudy (anonymous):

B. is correct .67 A. im not sure what i did wrong

OpenStudy (cggurumanjunath):

so u get 0.147 mol/l

OpenStudy (anonymous):

oh ok i didnt put the 7 after .14 so it told me it was wrong:/

OpenStudy (anonymous):

Thanks lots yaar!

OpenStudy (cggurumanjunath):

can u show steps for b ?

OpenStudy (anonymous):

yes. 32.7(1/56.37)=.580 Lsolution=857(1/1000)=.857 .580/.857 = .677M

OpenStudy (cggurumanjunath):

56.37 ?

OpenStudy (cggurumanjunath):

EXCELLENT !

OpenStudy (anonymous):

yes 56.37 molar mass of mgS Thank you, i know this for test bc of you:P

OpenStudy (cggurumanjunath):

:)